OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = x2 y = x + 2

OpenStudy (campbell_st):

Equate the equations sine both are equal to y \[x^2 = x + 2 \] or \[x^2 - x - 2= 0......(x -2)(x+1)=0 \] then x = 2 or -1 substitute the values into either equation y = 2^2 and y = (-1)^2 y = 4 and 1 ordered pairs are (2, 4) and (-1, 1)

OpenStudy (anonymous):

y = x + 2 We know Y =X2 So X2=X+2 X2-X-2=0 (X+1)(X-2)=0 X=-1 orX=2 So when X=-1----= Y=1 X=2---- = Y=4

OpenStudy (anonymous):

@campbell_st and samitheman99 Thank you so much.